# On the Entropy Calibration of Language Models

### v1

@@ 5.3 Main theorem
Why does future entropy preserve log loss? Future entropy adjustment can be derived as a first-order approximation of globally normalized temperature adjustment; we provide this derivation in Appendix B, along with a derivation in the MaxEnt RL framework (ziebart-2008-maximum). Global temperature adjustment attains calibration as long as the gradient of the log loss with respect to temperature is small (braverman2020calibration), which is a first-order condition. Then, intuitively, a first-order approximation of global temperature scaling should preserve this property.
The procedure described in Algorithm 1 is not practical to implement, as one would need to a fit a separate future entropy predictor for each generation step and each candidate token, each of which involves a slow data collection process based on a repeated sampling. Nonetheless, the existence of such an algorithm provides evidence that log loss tradeoffs are not inevitable in entropy calibration, despite the output space being exponentially large. One other point to note is that our analysis holds for any approximation of the future entropy that attains error $\delta$, with worse approximations just weakening the calibration error guarantee. For example, one could use the one-step future entropy (braverman2020calibration), or truncate to $k$ steps instead. We hope that our theory, which establishes future entropy as the target to approximate, guides future work to achieve better quality-diversity tradeoffs than existing approaches.
@@ 6 Additional Related Work
Calibration is most commonly studied in binary and multiclass classification, with some classic algorithms including binning, Platt scaling, and isotonic regression (PlattProbabilisticOutputs1999; zadrozny2002transforming; guo2017on; kumar2019verified). In the language modeling setting, liang2023holisticevaluationlanguagemodels evaluate the calibration of language models prompted to perform a wide range of classification tasks, finding that models are almost always miscalibrated and overconfident. In such a setting, one can simply apply standard calibration techniques to adjust the model’s outputted probabilities. More challenging is linguistic calibration, where models appear overconfident in the language they use to answer a question. To address this problem, past works propose techniques based on controllable generation and reinforcement learning (mielke-etal-2022-reducing; band2024linguistic). Finally, the term “calibration” is also used to describe the procedure of eliminating the model’s innate bias toward certain tokens when doing in-context learning, to improve task performance (pmlr-v139-zhao21c). All of these settings are distinct from our setting, which studies the calibration of a model’s entropy over an entire generation, and whose related work we discuss in Section 2.
@@ B.3 Future entropy adjustment from MaxEnt RL
The future entropy adjustment can also be derived in the MaxEnt RL framework (ziebart-2008-maximum), where the reward function is given by $r(x,y)=\log\hat{p}_{x}(y)$ with $\hat{p}$ denoting the base model. Specifically, we can write the MaxEnt RL objective as
@@ B.4 Scaling in the simplified setting
where $p_{i}=Z/i^{\alpha}$ is the power law probability of the $i$th item, with $Z=\sum_{i=1}^{v}1/i^{\alpha}$ denoting the normalizing constant. Next, taking $v\to\infty$ following the infinite urn setup in good-1953-population; karlin-1967-central, we compute
@@
as desired. This expression can also be found in Equation 23 of karlin-1967-central.
@@ Appendix C Experimental details
We study four model families (Qwen2.5 (0.5B, 1.5B, 3B, 7B, 14B, 32B, 72B) (qwen2025qwen25technicalreport), Llama 3 (1B, 3B, 8B, 70B) (grattafiori2024llama3herdmodels), Llama 2 (7B, 13B, 70B) (touvron2023llama2openfoundation), and Pythia (410M, 1.4B, 2.8B, 6.9B, 12B) (biderman2023pythiasuiteanalyzinglarge)) applied to the following three datasets:
@@
    WikiText-103 (merity2016pointersentinelmixturemodels): given 128 tokens of context from a Wikipedia passage, the model is tasked with completing the passage.
@@
    WritingPrompts (fan-etal-2018-hierarchical): given a writing prompt from the writingprompts subreddit along with 128 tokens of context from a human-written story, the model is tasked with completing the story.
@@
    CodeContests (li-2022-competition): given a coding problem from one of five websites (e.g. Codeforces) and 128 tokens of context from a human-written solution, the model is tasked with completing the solution.
In each setting, we use 5000 examples and limit the generation to at most 1024 tokens. For generation we use vLLM (kwon2023efficient) with the xFormers attention kernel (xFormers2022) and no quantization, and we use HuggingFace (wolf-etal-2020-transformers) with 4-bit quantization (dettmers2022bnb) to compute logprobs. All experiments are run using PyTorch (pytorch-2019), and all plots are produced using Matplotlib (Hunter:2007). For better readability, the entropy over time plots (Figure 3) are produced with exponential smoothing ($\alpha=0.2$). All experiments are run on 1-4 NVIDIA-A100-SXM4-80GB GPUs, or 1-4 NVIDIA RTX 6000 Ada Generation 49.1GB GPUs.

### v2

@@ 5.3 Main theorem
Why does future entropy preserve log loss? Future entropy adjustment can be derived as a first-order approximation of globally normalized temperature adjustment; we provide this derivation in Appendix B, along with a derivation in the MaxEnt RL framework (Ziebart et al., 2008). Global temperature adjustment attains calibration as long as the gradient of the log loss with respect to temperature is small (Braverman et al., 2020), which is a first-order condition. Then, intuitively, a first-order approximation of global temperature scaling should preserve this property.
The procedure described in Algorithm 1 is not practical to implement, as one would need to a fit a separate future entropy predictor for each generation step and each candidate token, each of which involves a slow data collection process based on a repeated sampling. Nonetheless, the existence of such an algorithm provides evidence that log loss tradeoffs are not inevitable in entropy calibration, despite the output space being exponentially large. One other point to note is that our analysis holds for any approximation of the future entropy that attains error $\delta$, with worse approximations just weakening the calibration error guarantee. For example, one could use the one-step future entropy (Braverman et al., 2020), or truncate to $k$ steps instead. We hope that our theory, which establishes future entropy as the target to approximate, guides future work to achieve better quality-diversity tradeoffs than existing approaches.
@@ 6 Additional Related Work
Calibration is most commonly studied in binary and multiclass classification, with some classic algorithms including binning, Platt scaling, and isotonic regression (Platt, 1999; Zadrozny and Elkan, 2002; Guo et al., 2017; Kumar et al., 2019). In the language modeling setting, Liang et al. (2023) evaluate the calibration of language models prompted to perform a wide range of classification tasks, finding that models are almost always miscalibrated and overconfident. In such a setting, one can simply apply standard calibration techniques to adjust the model’s outputted probabilities. More challenging is linguistic calibration, where models appear overconfident in the language they use to answer a question. To address this problem, past works propose techniques based on controllable generation and reinforcement learning (Mielke et al., 2022; Band et al., 2024). Finally, the term “calibration” is also used to describe the procedure of eliminating the model’s innate bias toward certain tokens when doing in-context learning, to improve task performance (Zhao et al., 2021). All of these settings are distinct from our setting, which studies the calibration of a model’s entropy over an entire generation, and whose related work we discuss in Section 2.
@@ B.3 Future entropy adjustment from MaxEnt RL
The future entropy adjustment can also be derived in the MaxEnt RL framework (Ziebart et al., 2008), where the reward function is given by $r(x,y)=\log\hat{p}_{x}(y)$ with $\hat{p}$ denoting the base model. Specifically, we can write the MaxEnt RL objective as
@@ B.4 Scaling in the simplified setting
where $p_{i}=Z/i^{\alpha}$ is the power law probability of the $i$th item, with $Z=\sum_{i=1}^{v}1/i^{\alpha}$ denoting the normalizing constant. Next, taking $v\to\infty$ following the infinite urn setup in Good (1953); Karlin (1967), we compute
@@
as desired. This expression can also be found in Equation 23 of Karlin (1967).
@@ Appendix C Experimental details
We study four model families (Qwen2.5 (0.5B, 1.5B, 3B, 7B, 14B, 32B, 72B) (Qwen et al., 2025), Llama 3 (1B, 3B, 8B, 70B) (Grattafiori et al., 2024), Llama 2 (7B, 13B, 70B) (Touvron et al., 2023), and Pythia (410M, 1.4B, 2.8B, 6.9B, 12B) (Biderman et al., 2023)) applied to the following three datasets:
@@
    WikiText-103 (Merity et al., 2017): given 128 tokens of context from a Wikipedia passage, the model is tasked with completing the passage.
@@
    WritingPrompts (Fan et al., 2018): given a writing prompt from the writingprompts subreddit along with 128 tokens of context from a human-written story, the model is tasked with completing the story.
@@
    CodeContests (Li et al., 2022): given a coding problem from one of five websites (e.g. Codeforces) and 128 tokens of context from a human-written solution, the model is tasked with completing the solution.
In each setting, we use 5000 examples and limit the generation to at most 1024 tokens. For generation we use vLLM (Kwon et al., 2023) with the xFormers attention kernel (Lefaudeux et al., 2022) and no quantization, and we use HuggingFace (Wolf et al., 2020) with 4-bit quantization (Dettmers et al., 2022) to compute logprobs. All experiments are run using PyTorch (Paszke et al., 2019), and all plots are produced using Matplotlib (Hunter, 2007). For better readability, the entropy over time plots (Figure 2) are produced with exponential smoothing ($\alpha=0.2$). All experiments are run on 1-4 NVIDIA-A100-SXM4-80GB GPUs, or 1-4 NVIDIA RTX 6000 Ada Generation 49.1GB GPUs.